Summary:

Overall a productive night. Photometric with OK seeing (0.9-1.5"). Part of first half lost to telescope and LBC troubles; the last 1.5 hours lost to wind. Did three pointings on UMaII; two photometric OBs of NGC galaxies and three pointings of Bootes.

Details

Observers: Xiaohui Fan, Linhua Jiang

Support: Olga Kuhn

Operator David Gonzalez Huerta

All Times UT

0000 beautiful clear sky. I will treat it as photometric. 25biases (only 23 in blue); 5biases

0120 left shutter door alarm.

0138 left shutter door finally opened.

0143 SkyFlatTest_Br.ob only about 300 counts per second.

0147 SkyFlat_Br_5.ob r: 30000 counts; B: 18000 counts. sky is getting dark quickly. only about 1/3 of the counts by the last image.

0204 G191B2B, spectrophotometric standard. focusing...

0246 struggled with focus for a while. maybe ok now. G191B2B_UsY.ob temp changing very quickly. standard star images look donuts. moving to SA 95

0309 SA95_Vr, SA95_BR, SA95_UI, SA95_UsY

0342 moving to SA101. doing a few things with the active optics.

0355 SA101_Vr, at airmass of 2.6. DX hits Z limit. stop and move on to UMaII

0412 focus seems OK, now starting UMII_pos0_Br.ob 6x300 sec in B and r.

0420 blue filter failed, and the blue image didn't read out. first red image is 1.5". wind is kind of high. restarting the OB.

0434 telescope oil pressure alarm; second image of the sequence garbage. on hold while fixing telescope. also rebooting camera.

0458 back on sky, do UMII_pos0_Br.ob again (from the start).

0515 telescope failed after second dither (so did blue controller temp.), blue image didn't read out. start again from the second dither positon.

0525 blue failed again... try one more time, if still doesn't work, we will switch to red only mode..

0610 doing the last four dithers of pos0; blue failed at dither position 4. let red to complete the sequence, then restart controller again... get blue one more chance. seeing for this sequence is about 1.3" in both B and r.

0623 after lbcfpia, go to UMII_pos1_Br.ob , 6x300 sec in B and r, seeing 1.3" - 1.5"

0709 after lbcfpia, UMII_pos2_Br.ob 6x300 sec in B and r, some blue images trailed; likely issue with slower blue guiding. seeing 1.1-1.3"

0754 going to SA98_UsUBVrRIY.ob. note that there is a lonely R standard lbcr.20080208.080541.fits which didn't have observer name set, so won't show in archive.

0808 going to SA101_UsUBVrRIY.ob.

0815 going to SA104_UsUBVrRIY.ob

0825 going to NGC_3184_photo. Another CCD temp failure, restart controller. 1x164 sec in (UV) and (RI) 1.0" seeing

0856 going to NGC_3351_photo 1x164 sec in (UV) and (RI) 1.0" seeing. blue still has enlongation issue. V image jumped (double peak)

0909 moving to Bootes field. adjusting SX...then focusing

0934 finally starting b23r.ob 5x240sec in U and Y, 1.0" seeing in Y, 1.5" in U for the 4 min exposure.

1001 focus then starting b30r.ob 5x240sec in U and Y. seeing 1.0 in Y, 1.5 in U.

1035 image looks good. keep going. b32r.ob 5x240sec in U and Y, seeing 0.9-1.1" in Y, 1.4-1.6" in U.

1102 focus then b34r.ob Us and Y. image 2 trailed, windy and have difficulty guiding. seeing is much worse, and wind is picking up.

1130 we need to close. Did 4 exposure on this, 2x240sec (one is useless) and 1x300sec. Probably need to re-do this OB

1220 40mph wind. time to call it a night...

-- XiaohuiFan - 07 Feb 2008
Topic revision: r7 - 08 Feb 2008, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback